# Biased AI can Influence Political Decision-Making

### v1

@@ 3.2 Biasing AI Language Model
Given that we do not want to inform participants of the bias of the models, this prefix is added to the participants’ input without their knowledge. However, using a prefix has its own challenges, specifically the models’ sensitivity to slight changes in prompts [sclar2023quantifying]. Therefore, we performed extensive testing of different prompts to elicit the most robust bias from our models. For more details on this experimentation, as well as the exact prefix’s used in our study, see Section A.4.
@@
Although the Political Compass Test results show that using a simple bias prefix is quite robust, we also include the exact opinions the models should have on each topic which match the given bias. For example, if we are asking a participant to learn about the topic “covenant marriage” and we want to induce a liberal bias, we would use the following prefix,
@@
See exact prompts in Section A.4.
@@ 3.3 Procedure
Before experimentation, participants were asked to sign an informed consent; however, the purpose of the study and any mention of biased AI were not included. Participants were only told they would be interacting with AI language models to complete tasks. Before the task portion of the experiment began, participants were asked demographic questions including their age, gender, race and ethnicity, their highest level of education, income, and political ideology. Then, participants were asked to complete two tasks, following a consistent three-stage design: initial choice section where their views on the topic were measured; interaction with a AI language model, where they gathered more information on the topic via typed conversation with the AI language model in a chatbox; and a post-choice section where they were again asked the same questions as the pre-choice section to measure how their opinions had changed. See Section A.1 for experimental overview.
We employed a experimental design, featuring three experimental factors (AI liberal bias, AI conservative bias, AI neutral) and two participant factors (Republican and Democrat participants). After consent and initial data gathering, participants were randomly assigned to an experimental condition (liberal bias AI, conservative bias AI, or neutral AI), an order of the tasks (Topic Opinion Task, and Budget Allocation Task), order of topics in the Topic Opinion Task (liberal support topic and conservative support topic), and specific topic for the Topic Opinion Task (select one of the two options per topic type in Table 3). Participants were not informed in any way as to whether the AI language model was biased or neutral. After completion of both tasks, we asked a series of follow-up question related to the participants’ experience with the AI language model and their overall level of AI knowledge, in general. Finally, we debriefed the participant on the true nature of the study, including the potential bias of the AI, and gave them an option to opt out of the study. No participant chose to opt out of the study.
@@ 3.4 Topic Opinion Task
In the Topic Opinion Task, participants were initially asked to express their opinions on various obscure political topics. We deliberately chose topics with clear political leanings but possessed a high degree of obscurity to minimize the likelihood that participants had strong opinions a priori. This was motivated by our desire to mitigate confirmation and implicit bias [taber_lodge_2006], as well as to model a real-world setting in which people would interact with AI to gain information on topics about which they know little. Although participants had little to no knowledge of these topics before interacting with the AI language model, the topics were chosen due to be those most likely to generate divided opinions based on political ideology in the U.S. (see Table 3). In the initial choice/opinion measurement, participants were given a 7-point Likert scaled question about how much they agreed or disagreed with a political statement, with a 0 indicating ‘I Don’t Know Enough to Say’.
@@ 3.6 Analysis
where are binary random variables indicating whether a participant was in the liberal or conservative bias experimental condition (if both are , this indicates a participant is in the control). Here represents the change in opinion as the difference between the response of post-opinion and pre-opinion questions, where . Note that the magnitude of represents the change in opinion, while the sign represents the direction of change (arbitrarily assigning negative to be more liberal and positive to be more conservative). We test the significance of and using an t-test with a significance threshold of .
@@
Again, note that we extend the model used in ANOVA when testing for the effects of prior knowledge and bias detection on the change in allocation ().
@@ 5 Author Contribution
The authors confirm contribution to the paper as follows: study conception and design: Jillian Fisher, Katharina Reinecke, Yulia Tsvetkov, Jennifer Pan, Daniel Fisher, Shangbin Feng, Yejin Choi; data collection: Jillian Fisher, Robert Aron; analysis and interpretation of results: Jillian Fisher, Thomas Richardson; draft manuscript preparation: Jillian Fisher, Jennifer Pan, Daniel Fisher, Katharina Reinecke, Yulia Tsvetkov. All authors reviewed the results and approved the final version of the manuscript.
@@ A.4.1 Prefix Selection Exploration
A.4.2 Selected Prefix
The process described in Section A.4.1 resulted in the following prefix’s being selected:
@@ B.4.1 Evaluate Persuasion Techniques
Due to the open nature of the Budget Allocation Task, we sought to determine if biased AI language models employed different persuasion techniques in their interactions with participants. To analyze the conversations, we used automatic annotation with GPT-4 [openai2023gpt4], employing detailed prompt engineering to identify various persuasion techniques in each Budget Allocation Task conversation. This annotation approach follows established practices in Natural Language Processing and has been shown to out-perform human annotation [chatgpt_annotates_well]. The list of persuasion techniques was derived from previous research [piskorski-etal-2023-semeval, zeng2024johnnypersuadellmsjailbreak], which itself was based on a meta-analysis of past studies. We note that only analysis from [piskorski-etal-2023-semeval] is shown in the main text, while the analysis using the list from [zeng2024johnnypersuadellmsjailbreak] can be found in Section E.4. We included two distinct lists to capture the breadth of persuasion techniques, which showed similar results. The full list of techniques is provided in the instructions below. We used the following instructions to guide the models annotations:
@@ E.1 Topic Opinion Task: Average Change in Opinion by Topic
However, this natural split is not seen in the liberal supported topics (bottom). We see that regardless of political partisanship of the participant, the average support under the control trends in support (positive). Interestingly enough, this is seen in both topics (Lacey Act of 1900 and Multifamily zoning). This means we had a ceiling effect when testing for statistical effects of the liberal bias AI, which might be one reason they resulted in non-significance.
@@
Participant Partisanship | Coefficient | Beta Value | t Value | p-value |
@@
Participant Partisanship | Coefficient | Value | t Value | p-value |
@@
E.2 Topic Opinion Task Results: No Prior Knowledge Subset
@@ E.3 AI Knowledge and Bias Detection Full Results
Participants | Coefficient | Beta Value | t-value | p-value |
@@
Participants | Coefficient | Beta Value | t-value | p-value |
@@
Participants | Coefficient | Beta Value | t-value | p-value |
@@
Participants | Coefficient | Beta Value | t-value | p-value |

### v2

@@ 3.2 Biasing the AI Language Model
Given that we do not want to inform participants of the bias of the models, this prefix is added to the participants’ input without their knowledge. However, using a prefix has its own challenges, specifically the models’ sensitivity to slight changes in prompts [sclar2023quantifying]. Therefore, we performed extensive testing of different prompts to elicit the most robust bias from our models. For more details on this experimentation, as well as the exact prefixes used in our study, see Section A.4.
@@
Although the Political Compass Test results show that using a simple bias prefix is quite robust, we also added the exact opinions the models should have on each topic which match the given bias in the prefix as well. For example, if we are asking a participant to learn about the topic “covenant marriage” and we want to induce a liberal bias, we would use the following prefix,
@@
We note that this prefix was appended to each input by the participant. See exact prompts in Section A.4.
@@ 3.3 Procedure
Before experimentation, participants were asked to sign an informed consent; however, the purpose of the study and any mention of biased AI were not included. Participants were only told they would be interacting with AI language models to complete tasks. Before the task portion of the experiment began, participants were asked demographic questions including their age, gender, race and ethnicity, their highest level of education, income, and political ideology. Then, participants were asked to complete two tasks, following a consistent three-stage design: an initial choice section where their views on the topic were measured; interaction with a AI language model, where they gathered more information on the topic via typed conversation with the AI language model in a chatbox; and a post-choice section where they were again asked the same questions as the pre-choice section to measure how their opinions had changed. See Section A.1 for experimental overview.
We employed a experimental design, featuring three experimental factors (AI liberal bias, AI conservative bias, AI neutral) and two participant factors (Republican and Democrat participants). After consent and initial data gathering, participants were randomly assigned to an experimental condition (liberal biased AI, conservative biased AI, or neutral AI), an order of the tasks (Topic Opinion Task, and Budget Allocation Task), order of topics in the Topic Opinion Task (liberal support topic and conservative support topic), and specific topic for the Topic Opinion Task (assign one of the two options per topic type in Table 3). Participants were not informed in any way as to whether the AI language model was biased or neutral. After completion of both tasks, we asked a series of follow-up questions related to the participants’ experience with the AI language model and their overall level of AI knowledge, in general. Finally, we debriefed the participant on the true nature of the study, including the potential bias of the AI, and gave them an option to opt out of the study. No participant chose to opt out of the study.
@@ 3.4 Topic Opinion Task
In the Topic Opinion Task, participants were initially asked to express their opinions on various obscure political topics. We deliberately chose topics with clear political leanings but also possessed a high degree of obscurity to minimize the likelihood that participants had strong opinions a priori. This was motivated by our desire to mitigate confirmation and implicit bias [taber_lodge_2006], as well as to model a real-world setting in which people would interact with AI to gain information on topics about which they know little. Although participants had little to no knowledge of these topics before interacting with the AI language model, the topics were chosen due to their divided opinions based on political ideology in the U.S. (see Table 3). In the initial choice/opinion measurement, participants were given a 7-point Likert scaled question about how much they agreed or disagreed with a political statement, with a 0 indicating ‘I Don’t Know Enough to Say’.
@@ 3.6 Analysis
where are binary random variables indicating whether a participant was in the liberal or conservative bias experimental condition (if both are , this indicates a participant is in the control). Here represents the change in opinion as the difference between the response of post-opinion and pre-opinion questions, where . Note that the magnitude of represents the change in opinion, while the sign represents the direction of change (arbitrarily assigning negative to be more liberal and positive to be more conservative). We test the significance of and using a t-test with a significance threshold of .
@@
Again, note that we extend the model used in the ANOVA when testing for the effects of prior knowledge and bias detection on the change in allocation ().
@@ 5 Author Contribution
The authors confirm their contribution to the paper as follows: study conception and design: Jillian Fisher, Katharina Reinecke, Yulia Tsvetkov, Jennifer Pan, Daniel Fisher, Shangbin Feng, Yejin Choi; data collection: Jillian Fisher, Robert Aron; analysis and interpretation of results: Jillian Fisher, Thomas Richardson; draft manuscript preparation: Jillian Fisher, Jennifer Pan, Daniel Fisher, Katharina Reinecke, Yulia Tsvetkov. All authors reviewed the results and approved the final version of the manuscript.
@@ A.4.1 Prefix Selection Exploration
A.4.2 Selected Prefixes
The process described in Section A.4.1 resulted in the following prefix’s being selected. We note that this prefix is added to every input from the user.
@@ B.4.1 Evaluate Persuasion Techniques
Due to the open nature of the Budget Allocation Task, we sought to determine if biased AI language models employed different persuasion techniques in their interactions with participants. To analyze the conversations, we used automatic annotation with GPT-4 [gpt4], employing detailed prompt engineering to identify various persuasion techniques in each Budget Allocation Task conversation. This annotation approach follows established practices in Natural Language Processing and has been shown to out-perform human annotation [chatgpt_annotates_well]. The list of persuasion techniques was derived from previous research [piskorski-etal-2023-semeval, zeng2024johnnypersuadellmsjailbreak], which itself was based on a meta-analysis of past studies. We note that only analysis from [piskorski-etal-2023-semeval] is shown in the main text, while the analysis using the list from [zeng2024johnnypersuadellmsjailbreak] can be found in Section E.4. We included two distinct lists to capture the breadth of persuasion techniques, which showed similar results. The full list of techniques is provided in the instructions below. We used the following instructions to guide the models annotations:
@@ E.1 Topic Opinion Task: Average Change in Opinion by Topic
However, this natural split is not seen in the liberal supported topics (bottom). We see that regardless of political partisanship of the participant, the average support under the control trends in support (positive). Interestingly enough, this is seen in both topics (Lacey Act of 1900 and Multifamily zoning). This means we had a ceiling effect when testing for statistical effects of the liberal biased AI, which might be one reason they resulted in non-significance.
@@
Participant Partisanship | Treatment Bias | Beta Value | t Value | p-value |
@@
Participant Partisanship | Treatment Bias | Value | t Value | p-value |
@@
E.2 Topic Opinion Task: No Prior Knowledge Subset
@@ E.3 AI Knowledge and Bias Detection Full Results
Participants | Treatment Bias | Beta Value | t-value | p-value |
@@
Participants | Treatment Bias | Beta Value | t-value | p-value |
@@
Participants | Treatment Bias | Beta Value | t-value | p-value |
@@
Participants | Treatment Bias | Beta Value | t-value | p-value |